OpenStudy (anonymous):

Solve 8^x = 4^(x + 1) a. x = -1 b. x = 2 c. x = 4 d. x = -2

OpenStudy (anonymous):

x = 2

OpenStudy (anonymous):

could you explain?

OpenStudy (anonymous):

rewrite 2^(3x)=2^(2x+2) take log_2 on both sides. 3x=2x+2 x=2